OpenStudy (anonymous):

How do I solve the quadratic equation y^2 - 7y - 18 = 0 by factoring

OpenStudy (anonymous):

@iits_avex be respectful buddy

OpenStudy (anonymous):

If you were to factor that, you need to find two products of the last term that add up to the middle term.

OpenStudy (anonymous):

Is it the same as completing a square?

OpenStudy (anonymous):

Hey man, what's your deal? This place is for asking/answering questions so that we can help one another. if you wanna pick on somebody or just be an idiot, go do it someplace else.

OpenStudy (anonymous):

Is it the same process as completing a square?

OpenStudy (anonymous):

Not necessarily. What two factors of -18 add up to -7?

OpenStudy (anonymous):

-9 and 2

OpenStudy (anonymous):

That is not a choice though... do the signs change at any point in the process?

OpenStudy (anonymous):

Ok. Now all you do is fill that into your parenthesis. your answer is (y-9)(y+2)

OpenStudy (anonymous):

That doesn't help because it isn't a choice... These are the choices: {-2, 9} {-9, -2} {-3, -6} {-6, 3}

OpenStudy (mathmale):

The poster of these crude messages has been suspended. Sorry for the interference with your discussion.

OpenStudy (anonymous):

@HackedSchool first option is an answer as explained by @rgonzalez (y-9) (y+2) = 0 now we will write y-9 =0 OR y+2 = 0 y =9 OR y= -2

OpenStudy (anonymous):

I see it now! Thanks, @rgonzalez and @niksva